OpenStudy (4meisu):

c. 30.0mL of 0.20M HCl will be neutralized with 0.50M NaOH. What volume of the base is needed?

OpenStudy (anonymous):

\[30 ml HCl (\frac{.20 mole HCl}{1000ml HCl})(\frac{1 mole NAOH}{1 mole HCl})(\frac{1000 ml NAOH}{.5 mole NAOH})=\] Just multiply across and you get the answer.